1.     Suppose the market demand curve for a good is Qd=1000-10p and the market supply curve is given by Qs=100+20p required.

       i.           Calculate the equilibrium price.


     ii.           Calculate the equilibrium quantity.



   iii.           Illustrate the concept of market equilibrium



1
Expert's answer
2021-03-04T10:40:34-0500

(i) The market equilibrium occurs when "Q_d=Q_s." Let's equate market demand and supply functions and find the equilibrium price:


"1000-10P_E=100+20P_E,""900=30P_E,""P_E=\$30."

(ii) Let's substitute "P_E" into the market demand function and find the equilibrium quantity:


"Q_E=1000-10\cdot\$30=700."

(iii) Let's plot the market demand and supply curves in a graph and illustrate the concept of market equilibrium:



As we can see from the graph, the market equilibrium occurs when the supply and demand curves intersects (where the quantity demanded and quantity supplied are equal). The equilibrium price equals $30 and equilibrium quantity equals 700 units.